# Convolution estimates for the Boltzmann gain operator with hard-spheres

### v3

@@ 4.1. The $L_{v}^{\infty}$– estimate for $Q_{0}$
$Q_{0}^{sm}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}|f(v^{*})g(v_{1}^{*})|\mathds{1}_{|v|\leq 10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1},$
$Q_{0}^{r}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}|f(v^{*})g(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{|v_{1}^{*}|\geq|v|/10}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1},$
$Q_{0}^{s}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}|f(v^{*})g(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{|v_{1}^{*}|<|v|/10}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}.$
Estimate for $Q_{0}^{sm}$
$\frac{|u|^{\gamma}{\langle}v{\rangle}^{k}}{{\langle}v^{*}{\rangle}^{k}{\langle}v_{1}^{*}{\rangle}^{l}}\mathds{1}_{|v|\leq 10}\mathds{1}_{|v^{*}|^{2}\geq E/2}\lesssim E^{1/2}(1+E)^{-k/2}\lesssim 1,$
since $k>2$. Using the above estimate, we obtain
${\langle}v{\rangle}^{k}|Q_{0}^{sm}(f,g)(v)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{k}(v^{*})g_{l}(v_{1}^{*})|\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}=Q_{M}^{+}(|f_{k}|,|g_{l}|),$
where we recall the Maxwell operator notation from (3.1). Then, Theorem 3.1 implies
(4.6) $\|{\langle}v{\rangle}^{k}Q_{0}^{sm}(f,g)\|_{L_{v}^{\infty}}\lesssim\|f_{k}\|_{L_{v}^{p_{0}}}\|g_{l}\|_{L_{v}^{q_{0}}},$
since $p_{0}\leq 2$.
Estimate for $Q_{0}^{r}$
$\frac{|u|^{\gamma}{\langle}v{\rangle}^{k}}{{\langle}v^{*}{\rangle}^{k}{\langle}v_{1}^{*}{\rangle}^{l}}\mathds{1}_{|v_{1}^{*}|\geq|v|/10}\mathds{1}_{|v^{*}|>E/2}\lesssim{\langle}v{\rangle}^{k-l}E^{1/2}(1+E)^{-k/2}\lesssim(1+E)^{\frac{1-l}{2}},$
since $k\geq l>2$. Arguing as in the previous case, we obtain
(4.7) $\|{\langle}v{\rangle}^{k}Q_{0}^{r}(f,g)\|_{L_{v}^{\infty}}\lesssim\|f_{k}\|_{L_{v}^{p_{0}}}\|g_{l}\|_{L_{v}^{q_{0}}},$
since $p_{0}\leq 2$.
We now estimate the singular operator $Q_{0}^{s}$.
Estimate for $Q_{0}^{s}$
Using the inequality $|u|\lesssim E^{1/2}$ and the fact that $0<\gamma\leq 1$, we estimate
$\frac{|u|^{\gamma}{\langle}v{\rangle}^{k}}{{\langle}v^{*}{\rangle}^{k}{\langle}v_{1}^{*}{\rangle}^{l}}\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|>E/2}\mathds{1}_{|v_{1}^{*}|<|v|/10}\lesssim\frac{{\langle}v{\rangle}^{3}}{{\langle}v^{*}{\rangle}^{2}{\langle}v_{1}^{*}{\rangle}^{l}}\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|>E/2}\mathds{1}_{|v_{1}^{*}|<|v|/10}$
$\lesssim\frac{{\langle}v{\rangle}^{3}}{{\langle}u{\rangle}^{2}{\langle}v_{1}^{*}{\rangle}^{l}}\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|>E/2}\mathds{1}_{|v_{1}^{*}|<|v|/10}\leq\frac{{\langle}v{\rangle}^{3}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}R_{\sigma}^{-}(u){\rangle}^{2}{\langle}v_{1}^{*}{\rangle}^{l}}\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|>E/2}\mathds{1}_{|v_{1}^{*}|<|v|/10},$
where for the last inequality we used (2.11). However, when $|v_{1}^{*}|<|v|/10$, (2.7) and the triangle inequality imply
$|R_{\sigma}^{-}(u)|=|v-v_{1}^{*}|\geq|v|-|v_{1}^{*}|\geq\frac{9|v|}{10},$
therefore
(4.8) $\frac{|u|^{\gamma}{\langle}v{\rangle}^{k}}{{\langle}v^{*}{\rangle}^{k}{\langle}v_{1}^{*}{\rangle}^{l}}\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|>E/2}\mathds{1}_{|v_{1}^{*}|<|v|/10}\lesssim\frac{{\langle}v{\rangle}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v_{1}^{*}{\rangle}^{l}}\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|^{2}\geq E/2}.$
We take different cases for $1<p_{0}\leq 2$ and $p_{0}=1$.
$\bullet$ Case $1<p_{0}\leq 2$ : Using (4.8) and Hölder’s inequality, we obtain
@@
${\langle}v{\rangle}^{k}|Q_{0}^{s}(f,g)(v)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v_{1}^{*}{\rangle}^{l}}|f_{k}(v^{*})g_{l}(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|^{2}\geq E/2}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
@@
$=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\left(\frac{{\langle}v{\rangle}|f_{k}(v^{*})|}{{\langle}v_{1}^{*}{\rangle}^{l}}\right)\left(|g_{l}(v_{1}^{*})||\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}\right)\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|^{2}\geq E/2}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
(4.9) $\leq I_{1}^{1/p_{0}}(v)I_{2}^{1/q_{0}}(v),$
@@
$I_{1}(v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{p_{0}}|f_{k}(v^{*})|^{p_{0}}}{{\langle}v_{1}^{*}{\rangle}^{lp_{0}}}\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|^{2}\geq E/2}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1},$
$I_{2}(v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|g_{l}(v_{1}^{*})|^{q_{0}}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2q_{0}}\,d\sigma\,dv_{1}.$
Since $l>2$ and $p_{0}\leq 2$, (2.25) implies $I_{1}(v)\lesssim{\langle}v{\rangle}^{p_{0}-2}\|f_{k}\|_{L_{v}^{p_{0}}}^{p_{0}}\leq\|f_{k}\|_{L_{v}^{p_{0}}}^{p_{0}}$. By (2.19), we also have $I_{2}(v)\lesssim\|g_{l}\|_{L^{q_{0}}}^{q_{0}}$.
Using (4.9) and the estimates for $I_{1},I_{2}$, we obtain
(4.10) ${\langle}v{\rangle}^{k}|Q_{0}^{s}(f,g)|\lesssim\|f_{k}\|_{L_{v}^{p_{0}}}\|g_{l}\|_{L_{v}^{q_{0}}},\quad 1<p_{0}\leq 2.$
$\bullet$ Case $p_{0}=1$ : We use again (4.8) to bound
${\langle}v{\rangle}^{k}|Q_{0}^{s}(f,g)|$ $\lesssim\|g_{l}\|_{L_{v}^{\infty}}\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}|f_{k}(v^{*})|}{{\langle}v_{1}^{*}{\rangle}^{l}}\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|^{2}\geq E/2}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
(4.11) $\lesssim\|f_{k}\|_{L_{v}^{1}}\|g_{l}\|_{L_{v}^{\infty}},$
where for the last inequality we used again the fact that $l>2$ and (2.25).
By (4.10), (4.11), we conclude
(4.12) $\|{\langle}v{\rangle}^{k}Q_{0}^{s}(f,g)\|_{L_{v}^{\infty}}\lesssim\|f_{k}\|_{L_{v}^{p_{0}}}\|g_{l}\|_{L_{v}^{q_{0}}},\quad 1\leq p_{0}\leq 2.$
Conclusion of the proof
Combining (4.6), (4.7), (4.12), estimate (4.4) follows. ∎
4.2. The $L_{v}^{\infty}$– estimate for $Q_{1}$
Proposition 4.2.
Let $0<\gamma\leq 1$. Consider $1<p_{1},q_{1}<\infty$ with $p_{1}<2$ and $1\leq p_{2},q_{2}\leq\infty$ with $p_{2}<2$, such that $\frac{1}{p_{1}}+\frac{1}{q_{1}}=\frac{1}{p_{2}}+\frac{1}{q_{2}}=1$. Then for any $k\geq l>2$, there holds the estimate
(4.13) $\|{\langle}v{\rangle}^{k}Q_{1}(f,g)\|_{L_{v}^{\infty}}\lesssim\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{p_{1}}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{q_{1}}}+\|{\langle}v{\rangle}^{l}f\|_{L_{v}^{q_{2}}}\|{\langle}v{\rangle}^{k}g\|_{L_{v}^{p_{2}}},\quad\forall\,f,g\in C_{c}^{\infty}({\mathbb{R}}^{3}).$
Proof.
Let $k\geq l>2$ and fix $f,g\in C_{c}^{\infty}({\mathbb{R}}^{3})$. Denote $f_{l}:={\langle}v{\rangle}^{l}f$, $g_{k}:={\langle}v{\rangle}^{k}g$. We note that since $p_{1}<2$ and $p_{2}<2$, we have $p_{1}<2<q_{1}$ and $p_{2}<2<q_{2}$ respectively.
We use the triangle inequality to bound $Q_{1}$ as follows
(4.14) $|Q_{1}(f,g)|\lesssim|Q_{1}^{sm}(f,g)|+|Q_{1}^{r}(f,g)|+|Q_{1}^{s_{1}}(f,g)|+|Q_{1}^{s_{2}}(f,g)|,$
where
$Q_{1}^{sm}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}|f(v^{*})g(v_{1}^{*})|\mathds{1}_{|v|\leq 10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1},$
$Q_{1}^{r}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}|f(v^{*})g(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{|v^{*}|\geq|v_{1}^{*}|/10}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1},$
$Q_{1}^{s_{1}}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}|f(v^{*})g(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1},$
$Q_{1}^{s_{2}}(f,g)$ $=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|u|^{\gamma}|f(v^{*})g(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1},$
and given $v\in{\mathbb{R}}^{3}$, we define the singular sets
(4.15) $\mathcal{B}_{1}(v)$ $=\left\{(v_{1},\sigma)\in{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}\,:\,|v^{*}|<\frac{|v_{1}^{*}|}{10},\,\,\,\,|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|\geq{\langle}v{\rangle}^{-1/2}\right\},$
(4.16) $\mathcal{B}_{2}(v)$ $=\left\{(v_{1},\sigma)\in{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}\,:\,|v^{*}|<\frac{|v_{1}^{*}|}{10},\,\,\,\,|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|<{\langle}v{\rangle}^{-1/2}\right\}.$
Estimate for $Q_{1}^{sm}$
$\frac{|u|^{\gamma}{\langle}v{\rangle}^{k}}{{\langle}v^{*}{\rangle}^{l}{\langle}v_{1}^{*}{\rangle}^{k}}\mathds{1}_{|v|\leq 10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\lesssim E^{1/2}(1+E)^{-k/2}\lesssim 1,$
since $k>2$. Using the above estimate, we obtain
${\langle}v{\rangle}^{k}|Q_{1}^{sm}(f,g)(v)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}=Q_{M}^{+}(|f_{l}|,|g_{k}|),$
Then, Theorem 3.1 implies
(4.17) $\|{\langle}v{\rangle}^{k}Q_{1}^{sm}(f,g)\|_{L_{v}^{\infty}}\lesssim\|f_{l}\|_{L_{v}^{p_{1}}}\|g_{l}\|_{L_{v}^{q_{1}}},$
since $p_{1}<2$.
Estimate for $Q_{1}^{r}$
$\frac{|u|^{\gamma}{\langle}v{\rangle}^{k}}{{\langle}v^{*}{\rangle}^{k}{\langle}v_{1}^{*}{\rangle}^{l}}\mathds{1}_{|v|>10}\mathds{1}_{|v^{*}|\geq|v_{1}^{*}|/10}\mathds{1}_{|v^{*}|>E/2}\lesssim(1+E)^{\frac{1-l}{2}}\lesssim 1,$
since $l>2$. Arguing as in the previous case, we obtain
(4.18) $\|{\langle}v{\rangle}^{k}Q_{1}^{r}(f,g)\|_{L_{v}^{\infty}}\lesssim\|f_{k}\|_{L_{v}^{p_{1}}}\|g_{l}\|_{L_{v}^{q_{1}}},$
since $p_{1}<2$.
We now estimate the singular operators $Q_{1}^{s_{1}}$ and $Q_{1}^{s_{2}}$.
Estimate for $Q_{1}^{s_{1}}$
Using the facts $|u|\leq|v|+|v_{1}|\lesssim E^{1/2}$ and $0<\gamma\leq 1$, we have
(4.19) $\frac{|u|^{\gamma}{\langle}v{\rangle}^{k}}{{\langle}v^{*}{\rangle}^{l}{\langle}v_{1}^{*}{\rangle}^{k}}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\lesssim\frac{|u|^{\gamma}{\langle}v{\rangle}^{k-1}}{{\langle}v_{1}^{*}{\rangle}^{k}}\frac{{\langle}v{\rangle}}{{\langle}v^{*}{\rangle}^{l}}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\lesssim\frac{{\langle}v{\rangle}}{{\langle}v^{*}{\rangle}^{l}}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}.$
We take different cases for $1<p_{2}<2$ and $p_{2}=1$.
$\bullet$ Case $1<p_{2}<2$ : Using (4.19) and Hölder’s inequality, we obtain
@@ 4.2. The $L_{v}^{\infty}$– estimate for $Q_{1}$
${\langle}v{\rangle}^{k}|Q_{1}^{s_{1}}(f,g)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}}{{\langle}v^{*}{\rangle}^{l}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
$\leq\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\left(\frac{|f_{l}(v^{*})|}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/p_{2}}}\right)\left(\frac{{\langle}v{\rangle}|g_{k}(v_{1}^{*})||\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/p_{2}}}{{\langle}v^{*}{\rangle}^{l}}\right)$
$\hskip 113.81102pt\times\mathds{1}_{|v|>10}\mathds{1}_{|v_{1}^{*}|^{2}\geq E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\chi(\widehat{u}\cdot\sigma)\,d\sigma\,dv_{1}$
(4.20) $\lesssim I_{1}^{1/q_{2}}(v)I_{2}^{1/p_{2}}(v),$
where
$I_{1}(v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|f_{l}(v^{*})|^{q_{2}}\chi(\widehat{u}\cdot\sigma)}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2q_{2}/p_{2}}}\mathds{1}_{|v|>10}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\,d\sigma\,dv_{1},$

### v4

@@ 4.1. The $L_{v}^{\infty}$– estimate
$I_{1}(v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{1-\frac{q}{p}}|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{q(k-\gamma)}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2q/p}}|f_{l}(v^{*})|^{q}\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\,d\sigma\,dv_{1},$
@@
$I_{2}(v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{2}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v^{*}{\rangle}^{lp}}|g_{k}(v_{1}^{*})|^{p}\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{1}(v)}\,d\sigma\,dv_{1}.$
Since $|v_{1}^{*}|^{2}>\frac{E}{2}>\frac{|v|^{2}}{2}>8$ and $l>2/p$, we use (2.26) to bound
$I_{2}(v)\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v_{1}^{*}{\rangle}^{2}|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2}}{{\langle}v^{*}{\rangle}^{lp}}|g_{k}(v_{1}^{*})|^{p}\mathds{1}_{|v_{1}^{*}|>2}\,d\sigma\,dv_{1}\lesssim\|g_{k}\|_{L^{p}_{v}}^{p}.$
In order to estimate $I_{1}$, we use the definition (4.3) of $\mathcal{B}_{1}(v)$, followed by (2.6), (2.13) and the substitution $y:=u=v-v_{1}$ to write
$I_{1}(v)$ $\leq{\langle}v{\rangle}^{1-\frac{q}{p}}\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-R_{\sigma}^{+}(u))|^{q}\mathds{1}_{|v|>4}\frac{|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{q(k-\gamma)}\mathds{1}_{\sqrt{1-|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{2}}\geq{\langle}v{\rangle}^{-1/2}}}{\left(1-|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{2}\right)^{q/p}}\,dv_{1}\,d\sigma$
$={\langle}v{\rangle}^{1-\frac{q}{p}}\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-R_{\sigma}^{+}(y))|^{q}\mathds{1}_{|v|>4}\frac{|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{q(k-\gamma)}\mathds{1}_{\sqrt{1-|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{2}}\geq{\langle}v{\rangle}^{-1/2}}}{\left(1-|\widehat{R}_{\sigma}^{+}(y)\cdot\sigma|^{2}\right)^{q/p}}\,dy\,d\sigma.$
Now, we use Proposition 2.1 to substitute $\nu:=R_{\sigma}^{+}(y)$ on the right-hand side of the above bound, as well as the fact that $q(k-\gamma)>p(k-\gamma)\geq 2$, to obtain
$I_{1}(v)$ $\lesssim{\langle}v{\rangle}^{1-\frac{q}{p}}\int_{{\mathbb{S}}^{2}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q}\mathds{1}_{|v|>4}\frac{|\widehat{\nu}\cdot\sigma|^{q(k-\gamma)}\mathds{1}_{\sqrt{1-|\widehat{\nu}\cdot\sigma|^{2}}\geq{\langle}v{\rangle}^{-1/2}}}{|\widehat{\nu}\cdot\sigma|^{2}\left(1-|\widehat{\nu}\cdot\sigma|^{2}\right)^{q/p}}\,d\nu\,d\sigma$
$\leq{\langle}v{\rangle}^{1-\frac{q}{p}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q}\mathds{1}_{|v|>4}\int_{{\mathbb{S}}^{2}}\frac{\mathds{1}_{\sqrt{1-|\widehat{\nu}\cdot\sigma|^{2}}\geq{\langle}v{\rangle}^{-1/2}}}{\left(1-|\widehat{\nu}\cdot\sigma|^{2}\right)^{q/p}}\,d\sigma\,d\nu$
$\approx{\langle}v{\rangle}^{1-\frac{q}{p}}\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q}\mathds{1}_{|v|>4}\int_{0}^{1}\frac{\mathds{1}_{\sqrt{1-x^{2}}\geq{\langle}v{\rangle}^{-1/2}}}{(1-x^{2})^{q/p}}\,dx\,d\nu.$
Now, the inner integral can be estimated as follows
$\int_{0}^{1}$ $\frac{\mathds{1}_{\sqrt{1-x^{2}}\geq{\langle}v{\rangle}^{-1/2}}}{(1-x^{2})^{q/p}}\,dx\lesssim 1+\int_{1/\sqrt{2}}^{1}\frac{\mathds{1}_{\sqrt{1-x^{2}}\geq{\langle}v{\rangle}^{-1/2}}}{(1-x^{2})^{q/p}}x\,dx\approx 1+\int_{{\langle}v{\rangle}^{-1}}^{1/2}z^{-q/p}\,dz\lesssim{\langle}v{\rangle}^{\frac{q}{p}-1},$
where we used the substitution $z:=1-x^{2}$, and the fact that $p<q$. We obtain
$I_{1}(v)\lesssim\int_{{\mathbb{R}}^{3}}|f_{l}(v-\nu)|^{q}\,d\nu=\|f_{l}\|_{L^{q}_{v}}^{q}.$
Combining (4.8) with the estimates for $I_{1},I_{2}$, we obtain
(4.9) $\|{\langle}v{\rangle}^{k+1-\gamma}Q_{s,1}^{+}(f,g)\|_{L_{v}^{\infty}}\lesssim\|f_{l}\|_{L_{v}^{q}}\|g_{k}\|_{L_{v}^{p}}.$
Estimate for $Q_{s,2}^{+}$
Using (4.7), the exponents’ compatibility condition, and Hölder’s inequality, we obtain
${\langle}v{\rangle}^{k+1-\gamma}|Q_{s,2}^{+}(f,g)|$ $\lesssim\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{k-\gamma}{\langle}v{\rangle}}{{\langle}v^{*}{\rangle}^{l}}|f_{l}(v^{*})g_{k}(v_{1}^{*})|\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\,d\sigma\,dv_{1}$
$=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\left(\frac{{\langle}v{\rangle}^{\frac{1}{p}-\frac{1}{q}}|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{k-\gamma}|f_{l}(v^{*})|}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/q}}\right)\left(\frac{{\langle}v{\rangle}^{2/q}|g_{k}(v_{1}^{*})||\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2/q}}{{\langle}v^{*}{\rangle}^{l}}\right)$
$\hskip 113.81102pt\times\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\,d\sigma\,dv_{1}$
(4.10) $\leq J_{1}^{1/p}(v)J_{2}^{1/q}(v),$
@@
$J_{1}(v)$ $:=\int_{{\mathbb{R}}^{3}\times{\mathbb{S}}^{2}}\frac{{\langle}v{\rangle}^{1-\frac{p}{q}}|\widehat{R}_{\sigma}^{+}(u)\cdot\sigma|^{p(k-\gamma)}}{|\widehat{R}_{\sigma}^{-}(u)\cdot\sigma|^{2p/q}}|f_{l}(v^{*})|^{p}\mathds{1}_{|v|>4}\mathds{1}_{|v_{1}^{*}|^{2}>E/2}\mathds{1}_{(v_{1},\sigma)\in\mathcal{B}_{2}(v)}\,d\sigma\,dv_{1},$